### v1

@@ 1 Introduction
In the past few years, modular approaches, utilizing separate architectures for understanding [2, 65, 3, 13, 14, 12], generation [40, 36, 59, 24, 34, 20], and editing tasks [85, 5, 32, 87, 46, 86], dominated multi-modal frameworks. While producing capable systems, it inherently creates complex, multi-stage pipelines. Such designs often suffer from architectural bottlenecks that limit deep, early-stage fusion of cross-modal information and introduce significant inference latency, presenting a major barrier to both efficiency and performance. In response to these limitations, the field has seen a rapid convergence towards unified multimodal modeling, aiming to integrate these disparate capabilities within a single, end-to-end architecture [67, 62, 75, 76, 11, 16]. Despite the trend towards unification, many current models remain tethered to the modular paradigm. We contend that a true paradigm shift—necessary to unlock the full potential of unified systems—demands a move towards a more fundamental, first-principles architecture that eschews external components. We therefore propose that a pure decoder-only transformer, trained under a unified objective, provides not just a sufficient, but a more elegant and potent foundation for the next generation of general-purpose multimodal intelligence.
@@ 2.2 Decoder-only MLLMs
Decoder-only MLLMs, also known as monolithic MLLMs, have recently emerged as a minimalist yet powerful alternative to the conventional compositional architecture. This paradigm aims to achieve greater efficiency and a more direct early fusion of modalities by eliminating the separate vision encoder. For example, Fuyu-8B [4] processes vision patches by feeding them through a simple linear patch embedding layer directly into the LLM backbone, which markedly reduce inference latency. Inspired by this success, subsequent works [18, 49, 19] further advance decode-only MLLMs by targeting their training processes and architectures. EvE [18] introduces a novel training objective that guides the model’s visual learning by aligning the last layer of LLM’s hidden states of image patch tokens with semantic features from a pre-trained vision encoder. Differently, Mono-InternVL [49] and EvEv2.0 [19] adopt a Mixture-of-Experts (MoE) framework, introducing a dedicated visual expert to handle visual-specific features more effectively. HoLVE [61] prepends a causal transformer to the LLM to explicitly convert both visual and textual inputs into a shared space. Despite these promising advancements, the overall training efficiency of these models remains a significant challenge. More importantly, the potential for the decoder-only architecture to create unified models that can seamlessly integrate multimodal understanding, generation, and even image editing capabilities remains a largely unexplored research avenue.
@@ 2.3 Unified Visual Understanding and Generation
Building on the success of MLLMs, the convergence of visual understanding and generation into a unified framework now represents a key research frontier. Pioneering unified MLLMs such as Chameleon [62],Transfusion [90], emu3 [67] and show-o [75] utilize visual tokenizers (e.g., VQ-VAE) to convert images into a sequence of discrete tokens, enabling seamless multimodal understanding and generation within a single model. However, the discretization inevitably results in lossy visual information and weakens in extracting semantic contents. Janus series [69, 11] decouples visual encoding for understanding and generation using two separate encoders, but may compromise performance due to task conflicts in shared LLM parameter space. Metaqueries [53], BLIP3-O [8], Uniworld-V [43] assembles off-the-shelf specialized MLMMs and diffusion models by tuning adapters and learnable query tokens, which sacrifices true architectural unification for modularity. BAGEL [16] and Mogao [41] employ a Mixture-of-Transformers (MoT) architecture, dedicating different components to autoregressive text generation and diffusion-based visual generation. However, while powerful, this hybrid approach inherits the significant inference latency of diffusion models and still requires separate encoders and tokenizers during the inference.
@@ 4 Model Training Pipeline
Figure 5: Overview of the training pipeline. In Stage 1, we first prepare a teacher model by training a two-layer MLP to connect InterViT [14] and the Qwen2.5 LLM [77]. This teacher model is then used to perform understanding distillation for the Und. FFN and the Patch Embedding layer. Simultaneously, we perform generation pretraining to optimize the Gen. FFN. All other parameters of the LLM remain frozen to preserve its pretrained language capabilities. In Stage 2 and 3, the entire model is unfrozen to conduct unified mid-training and supervised fine-tuning (SFT), respectively. The VAE component for visual generation is omitted from the figure for clarity.
@@ 4.1.1 Stage 1-1: Teacher Training
Rather than employing an off-the-shelf MLLM as teacher (e.g., Qwen2.5-VL [3]), we construct a custom teacher model to ensure parameter consistency between the LLM backbones of the teacher and student models, thereby improving distillation efficiency. Specifically, the teacher is built by connecting a pre-trained vision encoder (InterViT [14]) and a LLM (Qwen2.5 [77]) with a two-layer MLP as connector. During this sub-stage, we freeze the ViT and LLM, and exclusively train the MLP connector on a curated small-scale dataset of image-to-text caption pairs using the standard NTP loss.
@@ 6 Implementation details
Model Configurations: We conduct our experiments on two model variants, OneCAT-1.5B and OneCAT-3B. The OneCAT-1.5B model is based on Qwen2.5-1.5B-instruct [3] and contains 1.5B active parameters (4.5B total). The OneCAT-3B model is built upon Qwen2.5-3B-instruct [3] and utilizes 3B active parameters (9B total). For understanding distillation, the ViT of teacher model is InterViT-300M [14] for both two variants and the LLM of teacher model is Qwen2.5-1.5B-instruct and Qwen2.5-3B-instruct for OneCAT-1.5B and OneCAT-3B, respectively, to align with the LLM backbone of our OneCAT. Detailed instruction of the model configurations is shown in Tab. 2.
@@ 7 Evaluation
As shown in Tab. 3 and Tab. 4, we compare OneCAT with three types of models: encoder-based understanding-only models, encoder-free understanding-only models, and unified MLLMs. Our OneCAT-3B model demonstrates superior performance, significantly outperforming all existing encoder-free understanding-only MLLMs, e.g., HoVLE [61] and EvEv2 [19], across nearly all benchmarks. For instance, on OCR-related tasks including TextVQA (73.9), ChartQA (81.2), InfoVQA (64.8), and DocVQA (91.2), OneCAT-3B achieves new state-of-the-art results among encoder-free models. It also excels in general vision-language benchmarks such as MME-P (1630), MMBench (78.8), and MM-Vet (52.2).
@@ 7.3 Comparison of Inference Efficiency
Table 8: Efficiency comparison of OneCAT-3B and QwenVL-3B. Models are tested based on one NVIDIA H800 GPU. We report the time to the first token (TTFT) to measure the reduction of computational cost of prefilling phase. 256* denotes the number of visual tokens for thumbnail.

### v2

@@ 1 Introduction
In the past few years, modular approaches, utilizing separate architectures for understanding [2, 69, 3, 13, 14, 12], generation [42, 37, 61, 24, 34, 20], and editing tasks [90, 5, 32, 92, 48, 91], dominated multi-modal frameworks. While producing capable systems, it inherently creates complex, multi-stage pipelines. Such designs often suffer from architectural bottlenecks that limit deep, early-stage fusion of cross-modal information and introduce significant inference latency, presenting a major barrier to both efficiency and performance. In response to these limitations, the field has seen a rapid convergence towards unified multimodal modeling, aiming to integrate these disparate capabilities within a single, end-to-end architecture [71, 65, 80, 81, 11, 16]. Despite the trend towards unification, many current models remain tethered to the modular paradigm. We contend that a true paradigm shift—necessary to unlock the full potential of unified systems—demands a move towards a more fundamental, first-principles architecture that eschews external components. We therefore propose a pure decoder-only transformer, trained under a unified objective, providing not just a sufficient, but a more elegant and potent foundation for the next generation of general-purpose multimodal intelligence.
@@ 2.2 Decoder-only MLLMs
Decoder-only MLLMs, also known as monolithic MLLMs, have recently emerged as a minimalist yet powerful alternative to the conventional compositional architecture. This paradigm aims to achieve greater efficiency and a more direct early fusion of modalities by eliminating the separate vision encoder. For example, Fuyu-8B [4] processes vision patches by feeding them through a simple linear patch embedding layer directly into the LLM backbone, which markedly reduce inference latency. Inspired by this success, subsequent works [18, 51, 19, 35, 68, 62], further advance decode-only MLLMs by targeting their training processes and architectures. EvE [18] and VoRA [68] aligns the LLM’s hidden states with semantic features from a pre-trained ViT. However, directly using a smaller model (e.g., a ViT with hundreds of millions of parameters) as the teacher to distill knowledge into a significantly larger LLM (with several billion parameters) may restrict the expressive capacity of the LLM. Differently, Mono-InternVL [51] and EvEv2.0 [19] adopt a Mixture-of-Experts (MoE) framework, introducing a dedicated visual expert to handle visual-specific features more effectively. HoLVE [64] prepends a causal transformer to the LLM to explicitly convert both visual and textual inputs into a shared space. Despite these promising advancements, the overall training efficiency of these models remains a significant challenge. More importantly, the potential for the decoder-only architecture to create unified models that can seamlessly integrate multimodal understanding, generation, and even image editing capabilities remains a largely unexplored research avenue.
@@ 2.3 Unified Visual Understanding and Generation
Building on the success of MLLMs, the convergence of visual understanding and generation into a unified framework now represents a key research frontier [73, 11, 79, 95, 71, 80, 38, 55, 8, 45, 16, 43]. Pioneering unified MLLMs such as Chameleon [65],Transfusion [95], emu3 [71], show-o [80] and Synergen-vl [38] utilize visual tokenizers (e.g., VQ-VAE) to convert images into a sequence of discrete tokens, enabling seamless multimodal understanding and generation within a single model. However, the discretization inevitably results in lossy visual information and weakens in extracting semantic contents. Janus series [73, 11] decouples visual encoding for understanding and generation using two separate encoders, but may compromise performance due to task conflicts in shared LLM parameter space. Metaqueries [55], BLIP3-O [8], Uniworld-V [45] assembles off-the-shelf specialized MLMMs and diffusion models by tuning adapters and learnable query tokens, which sacrifices true architectural unification for modularity. BAGEL [16] and Mogao [43] employ a Mixture-of-Transformers (MoT) architecture, dedicating different components to autoregressive text generation and diffusion-based visual generation. However, while powerful, this hybrid approach inherits the significant inference latency of diffusion models and still requires separate encoders and tokenizers during the inference.
@@ 4 Model Training Pipeline
Figure 5: Overview of the training pipeline. In Stage 1, we first prepare a teacher model by training a two-layer MLP to connect InternViT [14] and the Qwen2.5 LLM [82]. This teacher model is then used to perform understanding distillation for the Und. FFN and the Patch Embedding layer. Simultaneously, we perform generation pretraining to optimize the Gen. FFN. All other parameters of the LLM remain frozen to preserve its pretrained language capabilities. In Stage 2 and 3, the entire model is unfrozen to conduct unified mid-training and supervised fine-tuning (SFT), respectively. The VAE component for visual generation is omitted from the figure for clarity.
@@ 4.1.1 Stage 1-1: Teacher Training
Rather than employing an off-the-shelf MLLM as teacher (e.g., Qwen2.5-VL [3]), we construct a custom teacher model to ensure parameter consistency between the LLM backbones of the teacher and student models, thereby improving distillation efficiency. Specifically, the teacher is built by connecting a pre-trained vision encoder (InternViT [14]) and a LLM (Qwen2.5 [82]) with a two-layer MLP as connector. During this sub-stage, we freeze the ViT and LLM, and exclusively train the MLP connector on a curated small-scale dataset of image-to-text caption pairs using the standard NTP loss.
@@ 6 Implementation details
Model Configurations: We conduct our experiments on two model variants, OneCAT-1.5B and OneCAT-3B. The OneCAT-1.5B model is based on Qwen2.5-1.5B-instruct [3] and contains 1.5B active parameters (4.5B total). The OneCAT-3B model is built upon Qwen2.5-3B-instruct [3] and utilizes 3B active parameters (9B total). For understanding distillation, the ViT of teacher model is InternViT-300M [14] for both two variants and the LLM of teacher model is Qwen2.5-1.5B-instruct and Qwen2.5-3B-instruct for OneCAT-1.5B and OneCAT-3B, respectively, to align with the LLM backbone of our OneCAT. Detailed instruction of the model configurations is shown in Tab. 2.
Table 3: Performance comparison across multiple multimodal understanding benchmarks. Higher scores are better, as indicated by the up-arrow ($\uparrow$). A-LLM denotes the number of activated LLM parameters, while Vis. indicates the parameter count of the vision encoder or tokenizer for multimodal understanding. Chameleon [65] does not report the parameter count of its vision tokenizer. slash (/) denotes that models do not require a vision encoder or tokenizer for multimodal understanding. Best in bold, second best is underlined (across unified models).
@@ 7 Evaluation
As shown in Tab. 3 and Tab. 4, we compare OneCAT with three types of models: encoder-based understanding-only models, encoder-free understanding-only models, and unified MLLMs. Our OneCAT-3B model demonstrates superior performance, significantly outperforming all existing encoder-free understanding-only MLLMs, e.g., HoVLE [64] and EvEv2 [19], across nearly all benchmarks. For instance, on OCR-related tasks including AI2D (77.8), ChartQA (81.2), InfoVQA (64.8), and DocVQA (91.2), OneCAT-3B achieves new state-of-the-art results among encoder-free models. It also excels in general vision-language benchmarks such as MME-P (1630), MMBench (78.8), and MM-Vet (52.2).
@@ 7.3 Comparison of Inference Efficiency
Table 8: Efficiency comparison of OneCAT-3B and Qwen2.5-VL-3B. Models are tested based on one NVIDIA H800 GPU. We report the time to the first token (TTFT) to measure the reduction of computational cost of prefilling phase. 256* denotes the number of visual tokens for thumbnail.